# Amortized Safe Active Learning for Real-Time Decision-Making: Pretrained Neural Policies from Simulated Nonparametric Functions

### v1

@@ Abstract
Active Learning (AL) is a sequential learning approach aiming at selecting the most informative data for model training. In many systems, safety constraints appear during data evaluation, requiring the development of safe AL methods. Key challenges of AL are the repeated model training and acquisition optimization required for data selection, which become particularly restrictive under safety constraints. This repeated effort often creates a bottleneck, especially in physical systems requiring real-time decision-making. In this paper, we propose a novel amortized safe AL framework. By leveraging a pretrained neural network policy, our method eliminates the need for repeated model training and acquisition optimization, achieving substantial speed improvements while maintaining competitive learning outcomes and safety awareness. The policy is trained entirely on synthetic data utilizing a novel safe AL objective. The resulting policy is highly versatile and adapts to a wide range of systems, as we demonstrate in our experiments. Furthermore, our framework is modular and we empirically show that we also achieve superior performance for unconstrained time-sensitive AL tasks if we omit the safety requirement.
Machine Learning, ICML
[…]
@@ 1 Introduction
Figure 1: The conventional safe AL relies on computationally expensive (orange) GP fitting and constrained acquisition optimization. Our amortized AL approach meta trains a safe active learner up-front from synthetic data, allowing a fast, easy and real-time applicable (green) deployment.
@@
Active learning (AL) is a sequential design of experiments, aiming to learn a task with reduced data labeling effort (Settles, 2010; Kumar & Gupta, 2020; Tharwat & Schenck, 2023). Each label is queried by optimizing an acquisition function, a function leveraging the current knowledge (typically model-based) to estimate the expected information gained from accessing new labels. AL is often discussed together with Bayesian optimization (BO), which aims to search global optima with limited evaluations (Srinivas et al., 2012; Brochu et al., 2010). The major difference is the acquisition function, where BO focuses only on candidate optima, while AL explores the complete space.
@@
In many engineering (Zimmer et al., 2018; Berkenkamp et al., 2016) or chemical design problems (Griffiths & Hernández-Lobato, 2020), data evaluations can trigger safety concerns. The safety constraints cannot be directly mapped to the input space, motivating the development of safe AL (Schreiter et al., 2015; Zimmer et al., 2018; Li et al., 2022) and safe BO (Sui et al., 2015; Yanan Sui et al., 2018; Berkenkamp et al., 2020). These approaches introduce additional model(s) to quantify the safety conditions and leverage the safety knowledge to constrain the acquisition optimization (Figure 1 left). In these problems, Gaussian processes (GPs, Rasmussen & Williams, 2006) are a popular tool. A GP is a distribution over functions, which provides well-calibrated predictive uncertainty estimates that are well suited for modeling safety confidence (Sui et al., 2015; Schreiter et al., 2015; Zimmer et al., 2018).
@@
Safe learning methods are prominent, capable of learning functions adaptively and safely, without requiring a parametric structure of the functions. However, the computation is heavy: (i) GPs scale cubically with the number of data points in time and are updated repeatedly; (ii) each data selection step solves an acquisition optimization, which can already be difficult for unconstrained learning (Swersky et al., 2020) and is even more complex under safety constraints (Sui et al., 2015; Schreiter et al., 2015). In particular, many physical systems require real-time decision-making (Nguyen–Tuong & Peters, 2010; Andersson et al., 2017; Lederer et al., 2021), which poses a significant challenge to employ these approaches. Various approximations have been investigated to make GP modeling more efficient (Titsias, 2009; Hensman et al., 2015; Bitzer et al., 2023), some of which have been adapted on scenarios of incremental data (Moss et al., 2023; Pescador-Barrios et al., 2024). However, and to the best of our knowledge, simplifying the repeated constrained acquisition optimization remains an open challenge.
@@
In this paper, we focus on safe AL applications (Schreiter et al., 2015; Zimmer et al., 2018; Li et al., 2022), and we wish to amortize the effort of the data querying process. Prior to the deployment, we propose to learn an AL policy based on synthetic nonparametric functions. The policy is a neural network (NN) which suggests each new query with a simple forward pass (Figure 1 right) hereby replacing both the GP modeling and the constrained optimization. Our approach also benefits unconstrained AL of GPs, by omitting the safety constraints. Note that we use the term model to refer to the model one wishes to actively learn on a specific task, while the NN policy guides the data collections and the data are then fed to model the task.
@@
In a nutshell, our approach (i) takes GPs as distributions of general nonparametric functions, (ii) utilizes a scalable random feature technique (Rahimi & Recht, 2007; Wilson et al., 2020) to generate functions in large scale, (iii) solves and meta learns AL decisions on those functions, and then (iv) zero-shot generalizes to real-world problems.
@@
  • We propose an NN policy for safe AL and, as an intermediate contribution, for unconstrained AL, that suggests new queries based on previously recorded data, hereby completely replacing the costly GP modeling and acquisition optimization steps during deployment;
  • We demonstrate that our policy leads to a tremendous speed gain, allowing for real-time applications, while still (i) showing competitive data quality for both unconstrained and safe AL, i.e. collected data lead to competitive modeling performance, and (ii) being safety aware when required;
  • Our policy is trained up-front exclusively on synthetic data, which represents a broad class of nonparametric functions, and generalizes to a multitude of systems;
  • In order to train our safe AL policy, we propose a differentiable safety-aware acquisition criterion in closed form, which serves as an objective function in our pretraining environment.
@@
Safe learning often employs GPs. Gelbart et al. (2014) proposed a constrained BO, discounting the acquisition scores by a GP safety confidence, assuming the acquisition function is non-negative. Schreiter et al. (2015); Sui et al. (2015) proposed to constrain the acquisition optimization by the GP safety confidence, leading to a learning scheme with probabilistic safety guarantees. This scheme was then extended to systems with multiple safety constraints (Berkenkamp et al., 2020), to time-series modeling (Zimmer et al., 2018), and to safe multi-task learning (Li et al., 2022) as well as transfer learning (Li et al., 2024). Another scheme of constrained sequential decision-making is safe reinforcement learning (García et al., 2015; Brunke et al., 2022; Gu et al., 2024). This is often based on a constrained Markov decision process, where an agent learns to make safe actions on states (given Markov property) (Altman, 1999). These safe learning methods employ significant computations, impeding their deployment in real-time decision-making problems.
Recently, various meta learning approaches have been investigated to streamline sequential learning methods. Rothfuss et al. (2021) proposed to meta learn GP hyperparameters from existing tasks, simplifying GP modeling in safe BO. Liu et al. (2020b); Bitzer et al. (2023) proposed to pretrain from synthetic data to infer the GP model parameters. On the acquisition phase, Swersky et al. (2020) proposed to correlate acquisition optimizations of different querying steps, achieving an amortized unconstrained BO. These methods either require existing meta tasks or do not consider constrained acquisition optimization, while our method simply queries end-to-end on constrained AL.
To further streamline the entire data selection cycle, Chen et al. (2017) proposed an NN optimizer for BO tasks, which bypassed modeling and optimization by directly inferring the next query from evaluated data. Foster et al. (2021); Ivanova et al. (2021) proposed the deep adaptive design (DAD), inferring the queries for an unconstrained AL of Bayesian parametric functions (which requires a known parametric structure, in contrast to our method). Chen et al. (2017); Foster et al. (2021); Ivanova et al. (2021) trained their NN policies purely with synthetic data. However, constrained data querying was not considered.
To the best of our knowledge, no literature has so far casted a constrained sequential learning task into an end-to-end learning approach as we propose in our work, i.e. learning to safely and adaptively learn.
@@ 2 Problem Statement and Assumptions
Algorithm 1 Safe AL with NN Policy
0:  $\mathcal{D}_{0}\subseteq\mathcal{X}\times\mathcal{Y}\times\mathcal{Z}$, AL policy $\phi$, $T$
1:  for $t=1,...,T$ do
2:     $\bm{x}_{t}=\phi(T-t+1,\mathcal{D}_{t-1})$
3:     Query at $\bm{x}_{t}$ to get $y_{t},z_{t}$
4:     $\mathcal{D}_{t}\leftarrow\mathcal{D}_{t-1}\cup\{\bm{x}_{t},y_{t},z_{t}\}$
5:  end for
6:  return $\mathcal{D}_{T}$ suitable to model $f$
@@
In conventional safe AL (Figure 1 left, Schreiter et al., 2015; Zimmer et al., 2018), each query at step $t+1$ is obtained by solving
$\begin{split}\text{argmax}_{\bm{x}\in\mathcal{X}}&a(\bm{x}|Y_{\text{init}},y_{1:t})\\ s.t.\ &p(z(\bm{x})\geq 0|Z_{\text{init}},z_{1:t})\geq 1-\gamma,\end{split}$ (1)
where $p(z(\bm{x})\geq 0|Z_{\text{init}},z_{1:t})$ is the predictive safety distribution, $a$ is an acquisition function, $\gamma\in[0,1]$ is a probability tolerance of being unsafe (usually $\gamma>0$), and the space of safe $\bm{x}$ is the safe set. The estimated safe set adapts to each new safety measurement (Sui et al., 2015; Yanan Sui et al., 2018). The acquisition function and safety distribution are computed from GPs. It is challenging to iteratively update GPs and solve the constrained optimization.
@@
In this paper, we aim to have a safe AL policy up-front, allowing us to replace the modeling and constrained acquisition optimization by simple policy forward passes (Figure 1 right and Algorithm 1). Formally speaking, the policy function takes as inputs a budget variable and an flexible size of observed data, and returns the next query proposal, i.e. $\phi:\mathbb{N}\times space(\mathcal{X}\times\mathcal{Y}\times\mathcal{Z})\rightarrow\mathcal{X}$. We will describe in Section 3.1.2 why a budget variable is included into the policy input. We assume no additional real data are available for the policy training.
Our paper investigates: (i) how to simulate safe AL tasks in large scale for our training, and (ii) how to design an effective training objective. We will first turn unconstrained AL into a learning problem, as inspired by Chen et al. (2017); Foster et al. (2021); Ivanova et al. (2021), and then further incorporate safety awareness. To achieve this, however, we need further assumptions on the functions $f$ and $q$.
@@
We assume the observation data are normalized to zero mean and unit variance. Furthermore, as existing safe learning methods (Sui et al., 2015; Schreiter et al., 2015), we assume that the nonparametric functions $f$ and $q$ have GP priors (Rasmussen & Williams, 2006). A GP is a distribution over functions, characterized by a mean (e.g., $\mathbb{E}[f]$) and a kernel function that specifies the covariance between function values at two input points, $\bm{x}$ and $\bm{x}^{\prime}$ (e.g., $\text{Cov}[f(\bm{x}),f(\bm{x}^{\prime})]$). A kernel function, typically parameterized, encodes the amplitude and smoothness of the functions. W.l.o.g., the prior mean is usually assumed zero, which holds true when the observation values are normalized. For the safety function, one could as well assume a zero mean prior, but a well-selected prior mean can sometimes be beneficial. For example, if the domain of interest $\mathcal{X}$ is chosen such that the center area is safe, then we may assign a prior mean which remains reasonably positive at the center but decreases to negative values at the boundary. The assumption is formally described below.
@@
The unknown function $f$ has a GP prior $\mathcal{GP}(\bm{0},k_{\theta})$, and likewise $q\sim\mathcal{GP}_{\theta_{q}}(\mu_{q},k_{q})$, where $k_{\theta}$ is parameterized by $\theta$, and $\mu_{q}$ and $k_{q}$ are jointly parameterized by $\theta_{q}$. The output and safety observations at $\bm{x}$ are characterized by $y(\bm{x})=f(\bm{x})+\epsilon$, $z(\bm{x})=q(\bm{x})+\epsilon_{q}$, where $\epsilon\sim\mathcal{N}(0,\sigma^{2}),\epsilon_{q}\sim\mathcal{N}(0,\sigma_{q}^{2})$ are i.i.d. Gaussian noises (Rasmussen & Williams, 2006). Here, $k_{\theta},k_{q}:\mathcal{X}\times\mathcal{X}\rightarrow\mathbb{R}$ are kernel functions, and $\mu_{q}:\mathcal{X}\rightarrow\mathbb{R}$ is the prior mean of $q$. We further assume $k_{\theta}(\bm{x},\bm{x}^{\prime}),k_{q}(\bm{x},\bm{x}^{\prime})\leq 1$.
Bounding the kernel scales by one is not restrictive, as we assume the observations are normalized to unit variance. Due to the GP priors, any finite number of functional values (or of the safety values) are jointly Gaussian. GP distributions are provided in closed form in Appendix A.
3 Pretrain Safe AL - NN Forward to Replace GP and Constrained Acquisition
Our goal here is to train a policy $\phi$ to run Algorithm 1. In other words, we construct our preparation block illustrated in Figure 1. Here we take inspiration from DAD (Foster et al., 2021; Ivanova et al., 2021). The idea is to exploit the GP priors (2.1) before safe AL experiments. We use $p(f),p(q)$ and the Gaussian likelihoods $p(y|\bm{x},f)=\mathcal{N}\left(y|f(\bm{x}),\sigma^{2}\right)$, $p(z|\bm{x},q)=\mathcal{N}\left(z|q(\bm{x}),\sigma_{q}^{2}\right)$ to construct a simulator. Notably, we go beyond DAD by considering the safety function $q$ and nonparametric priors on $f$ and $q$. This allows us to sample functions, simulate policy-based safe AL (Algorithm 1) and then meta optimize an objective function which encodes an acquisition criterion. Since we cannot know the GP hyperparameters for an unknown deployment task, we sample the hyperparameters to ensure that the policy learns a broad range of function classes.

### v2

# Amortized Safe Active Learning for Real-Time Data Acquisition: Pretrained Neural Policies from Simulated Nonparametric Functions

@@ Abstract
Safe active learning (AL) is a sequential scheme for learning unknown systems while respecting safety constraints during data acquisition. Existing methods often rely on Gaussian processes (GPs) to model the task and safety constraints, requiring repeated GP updates and constrained acquisition optimization—incurring in significant computations which are challenging for real-time decision-making. We propose an amortized safe AL framework that replaces expensive online computations with a pretrained neural policy. Inspired by recent advances in amortized Bayesian experimental design, we turn GPs into a pretraining simulator. We train our policy prior to the AL deployment on simulated nonparametric functions, using Fourier feature-based GP sampling and a differentiable, safety-aware acquisition objective. At deployment, our policy selects safe and informative queries via a single forward pass, eliminating the need for GP inference or constrained optimization. This leads to substantial speed improvements while preserving safety and learning quality. Our framework is modular and can be adapted to unconstrained, time-sensitive AL tasks by omitting the safety requirement.
@@ 1 Introduction
Figure 1: Conventional safe AL relies on computationally expensive (orange) GP fitting and constrained acquisition. Our amortized approach meta trains a safe learner up-front on synthetic data, allowing fast, real-time (green) deployment.
@@
Active learning (AL) is a sequential design of experiments, aiming to learn a task with reduced data labeling effort [43, 24, 49]. Each label is queried by optimizing an acquisition function, a function leveraging the current knowledge (typically model-based) to estimate the expected information gained from accessing new labels. AL is often discussed together with Bayesian optimization (BO), which aims to search global optima with limited evaluations [45, 8]. The major difference is the acquisition function, where BO focuses only on candidate optima, while AL explores the complete space.
@@
In many engineering [55, 3] or chemical design problems [15], data evaluations can trigger safety concerns. The safety constraints cannot be directly mapped to the input space, motivating the development of safe AL [40, 55, 26] and safe BO [46, 47, 4]. These approaches introduce additional model(s) to quantify safety conditions and constrain the acquisition optimization accordingly (Figure 1 left). Gaussian processes (GPs [37]) are widely used in this context due to their well-calibrated uncertainty estimates, which are well suited for modeling safety confidence.
@@
Safe learning methods are prominent, capable of learning functions adaptively and safely, without requiring a parametric structure of the functions. However, the computation is heavy: (i) GPs scale cubically with the number of data points and are updated repeatedly; (ii) each data selection step solves an acquisition optimization, which can already be difficult for unconstrained learning [48] and is even more complex under safety constraints [46, 40]. For instance, safety guarantees often rely on solving constrained acquisition via fine discretization [47, 4]. The expensive computation poses a particular challenge for physical systems that require real-time decision-making [31, 2, 25]. To alleviate this, efficient GP approximations have been investigated [50, 18, 6], including approaches designed for incremental data [30, 34].
@@
In this paper, we focus on safe AL problems [40, 55, 26], which are also relevant to safe BO, particularly in approaches that separate safe space exploration from BO phase [47, 7]. We aim to amortize the online data querying process. Inspired by recent amortized Bayesian experimental design literature [11, 20], we propose to learn an AL policy offline using synthetic functions. The policy is a neural network (NN) which suggests a new query via a simple forward pass (Figure 1 right) hereby replacing both the GP modeling and the constrained optimization at deployment. Our approach also benefits unconstrained AL, by omitting the safety constraints. Note that we use the term model to refer to the task-specific model being actively learned, while the NN policy guides the data collections, feeding the data to the model.
@@
In a nutshell, our approach (i) takes GPs as distributions of general nonparametric functions, (ii) utilizes a scalable Fourier feature technique [35, 53] to generate functions in large scale, (iii) solves and meta learns AL decisions on those functions, and then (iv) zero-shot generalizes to real-world problems.
@@
  • We propose an NN policy for safe AL and, as an intermediate contribution, for unconstrained AL, that suggests new queries based on previously recorded data, hereby completely replacing the costly GP modeling and acquisition optimization at deployment;
  • This leads to a tremendous speed gain, allowing for real-time data acquisition with modeling performance comparable to traditional AL methods, while maintaining safety when required;
  • Our policy is trained up-front exclusively on synthetic data of a broad class of functions sampled via GP Fourier features, enabling generalization across systems;
  • We introduce a closed-form, differentiable, safety-aware acquisition objective to train our safe AL policy, which may itself be also a novel acquisition criterion.
@@
Safe learning often employs GPs. Gelbart et al. [13] proposed constrained BO, discounting acquisition scores by a GP constraint confidence. Schreiter et al. [40], Sui et al. [46] directly constrained the acquisition optimization, leading to safe AL and safe BO with probabilistic safety guarantees. Sui et al. [47] introduced a stagewise safe BO that separates safe space exploration from BO phase, enabling AL methods to be applied specifically for safety [7]. Safe AL and safe BO were extended to systems with multiple safety constraints [4], time-series modeling [55], multi-task learning [26] and transfer learning [27]. High-dimensional safe BO has also been explored via approximations and multi-stage procedures [21, 7]. Safe learning methods employ significant computations, impeding their deployment in real-time problems.
Recently, meta-learning has been explored to streamline sequential learning. Rothfuss et al. [39] meta-learned GP priors from existing tasks, simplifying GP modeling in safe BO. Liu et al. [29], Bitzer et al. [6] pretrained on synthetic data to infer the GP parameters. On the acquisition phase, Swersky et al. [48] amortized unconstrained BO by correlating acquisition optimizations across querying steps. These methods either require existing meta tasks or do not consider constrained acquisition optimization, while our method simply queries end-to-end on constrained AL.
To further streamline the entire data selection cycle, Chen et al. [10] proposed an NN optimizer for BO tasks, which bypassed modeling and optimization by directly inferring the next query from evaluated data. Foster et al. [11], Ivanova et al. [20] proposed the deep adaptive design (DAD), inferring the queries for an unconstrained AL of Bayesian parametric functions (which requires a known parametric structure, in contrast to our method). Huang et al. [19] extended such amortized Bayesian experimental design to account for model-based decision-making. These approaches train NN policies purely on synthetic data. However, constrained data querying was not considered.
To the best of our knowledge, a constrained sequential learning has only been formulated as a learning problem in safe reinforcement learning [12, 9, 16], often based on a constrained Markov decision process where an agent learns to make safe actions on states (given the Markov property) [1]. In safe AL problems, where the Markov property is generally not fulfilled, no literature has explored an end-to-end learning approach as we propose in this work.
@@ 2 Problem statement and assumptions
Algorithm 1 Safe with NN Policy
1:$\mathcal{D}_{0}\subseteq\mathcal{X}\times\mathcal{Y}\times\mathcal{Z}$, AL policy $\phi$, $T$
2:for $t=1,...,T$ do
3:     $\bm{x}_{t}=\phi(T-t+1,\mathcal{D}_{t-1})$
4:     Query at $\bm{x}_{t}$ to get $y_{t},z_{t}$
5:     $\mathcal{D}_{t}\leftarrow\mathcal{D}_{t-1}\cup\{\bm{x}_{t},y_{t},z_{t}\}$
6:end for
7:return $\mathcal{D}_{T}$ suitable to model $f$
@@
Conventional safe AL (Figure 1 left and [40, 55]) obtains each query at step $t+1$ by solving
$\begin{split}\text{argmax}_{\bm{x}\in\mathcal{X}}a(\bm{x}|Y_{\text{init}},y_{1:t})\ s.t.\ p(z(\bm{x})\geq 0|Z_{\text{init}},z_{1:t})\geq 1-\gamma,\end{split}$ (1)
where $p(z(\bm{x})\geq 0|Z_{\text{init}},z_{1:t})$ is the predictive safety distribution, $a$ is an acquisition function, $\gamma\in[0,1]$ is a probability tolerance of being unsafe (usually $\gamma>0$), and the space of safe $\bm{x}$ is the safe set. The estimated safe set adapts to each new safety measurement [46, 47]. The acquisition function and safety distribution are computed from GPs. It is expensive to iteratively update GPs and solve the constrained optimization.
@@
In this paper, we aim to have a safe AL policy up-front, allowing us to replace the modeling and constrained acquisition optimization by simple policy forward passes (Figure 1 right and Algorithm 1). Formally speaking, the policy function takes as inputs a budget variable and a flexible size of observed data, and returns the next query proposal, i.e. $\phi:\mathbb{N}\times space(\mathcal{X}\times\mathcal{Y}\times\mathcal{Z})\rightarrow\mathcal{X}$. We will describe in Section 3.1.2 why a budget variable is included into the policy input. We assume no additional real data are available for the policy training.
Our paper investigates: (i) how to simulate safe AL tasks in large scale for our training, and (ii) how to design an effective training objective. We will first turn unconstrained AL into a learning problem, as inspired by Chen et al. [10], Foster et al. [11], Ivanova et al. [20], and then incorporate safety awareness. To achieve this, however, we need further assumptions on the functions $f$ and $q$.
@@
We assume the tasks are normalized to zero mean and unit variance. Furthermore, as existing safe learning methods (e.g. [55, 4]), we assume the unknown functions $f$ and $q$ have GP priors [37]. A GP is a distribution over functions, characterized by a mean (e.g., $\mathbb{E}[f]$) and a kernel specifying covariance between function values at two inputs, $\bm{x}$ and $\bm{x}^{\prime}$ (e.g., $\text{Cov}[f(\bm{x}),f(\bm{x}^{\prime})]$). A kernel, typically parameterized, encodes the function amplitude and smoothness. W.l.o.g., the prior mean is usually assumed zero, which holds true when the observation values are normalized. For the safety function, one could as well assume a zero mean prior, but a well-selected prior mean can sometimes be beneficial. For example, if the domain of interest $\mathcal{X}$ is chosen such that the center area is safe, then we may assign a prior mean which remains reasonably positive at the center but decreases to negative values at the boundary. Given GP priors, any finite number of functional values (or of the safety values) are jointly Gaussian. The assumption is formulated below, while closed-form GP distributions are detailed in Appendix A.
@@
The unknown functions follow GP priors: $f\sim\mathcal{GP}(\bm{0},k_{\theta}),q\sim\mathcal{GP}_{\theta_{q}}(\mu_{q},k_{q})$. $k_{\theta},k_{q}:\mathcal{X}\times\mathcal{X}\rightarrow\mathbb{R}$ are kernels, and $\mu_{q}:\mathcal{X}\rightarrow\mathbb{R}$ is a mean. $k_{\theta}$ is parameterized by $\theta$, and $\mu_{q}$ and $k_{q}$ are jointly parameterized by $\theta_{q}$. The output and safety observations at $\bm{x}$ are $y(\bm{x})=f(\bm{x})+\epsilon$, $z(\bm{x})=q(\bm{x})+\epsilon_{q}$, where $\epsilon\sim\mathcal{N}(0,\sigma^{2}),\epsilon_{q}\sim\mathcal{N}(0,\sigma_{q}^{2})$ are i.i.d. Gaussian noises [37]. We assume $k_{\theta}(\bm{x},\bm{x}^{\prime}),k_{q}(\bm{x},\bm{x}^{\prime})\leq 1$, as the data are assumed normalized.
3 Pretrain safe AL - NN to replace GP and constrained acquisition
Our goal here is to train a policy $\phi$ to run Algorithm 1. In other words, we construct our preparation block illustrated in Figure 1. Here we take inspiration from DAD [11, 20]. The idea is to exploit the GP priors (2.1) before safe AL experiments. We use $p(f),p(q)$ and the Gaussian likelihoods $p(y|\bm{x},f)=\mathcal{N}\left(y|f(\bm{x}),\sigma^{2}\right)$, $p(z|\bm{x},q)=\mathcal{N}\left(z|q(\bm{x}),\sigma_{q}^{2}\right)$ to construct a simulator. Notably, we go beyond DAD by considering the safety function $q$ and nonparametric priors on $f$ and $q$. This allows us to sample functions, simulate policy-based safe AL (Algorithm 1) and then meta optimize an objective function which encodes an acquisition criterion. Since we cannot know the GP hyperparameters for an unknown deployment task, we sample the hyperparameters to ensure that the policy learns a broad range of function classes.